# SemNav: A Semantic Segmentation-Driven Approach to Visual Semantic Navigation

### v1

@@ Abstract
Visual Semantic Navigation (VSN) is a fundamental problem in robotics, where an agent must navigate toward a target object in an unknown environment, mainly using visual information. Most state-of-the-art VSN models are trained in simulation environments, where rendered scenes of the real world are used, at best. These approaches typically rely on raw RGB data from the virtual scenes, which limits their ability to generalize to real-world environments due to domain adaptation issues. To tackle this problem, in this work, we propose SemNav, a novel approach that leverages semantic segmentation as the main visual input representation of the environment to enhance the agent’s perception and decision-making capabilities. By explicitly incorporating high-level semantic information, our model learns robust navigation policies that improve generalization across unseen environments, both in simulated and real world settings. We also introduce a newly curated dataset, i.e. the SemNav dataset, designed for training semantic segmentation-aware navigation models like SemNav. Our approach is evaluated extensively in both simulated environments and with real-world robotic platforms. Experimental results demonstrate that SemNav outperforms existing state-of-the-art VSN models, achieving higher success rates in the Habitat 2.0 simulation environment, using the HM3D dataset. Furthermore, our real-world experiments highlight the effectiveness of semantic segmentation in mitigating the sim-to-real gap, making our model a promising solution for practical VSN-based robotic applications. We release SemNav dataset, code and trained models at: https://github.com/gramuah/semnav
@@ I Introduction
Autonomous navigation remains a fundamental challenge in robotics, particularly in unstructured and dynamic environments. Traditional approaches, such as the ones using Simultaneous Localization and Mapping (SLAM), for example, focus on a geometric reconstruction of the environment in conjunction with path planning and obstacle avoidance techniques (e.g. [1, 2, 3]). However, these classical models typically struggle when they need to generalize across environments and domains [4].
Visual Semantic Navigation (VSN) has emerged as an alternative paradigm that leverages advances in machine learning and the availability of large-scale data. VSN models are essentially learning-based navigation approaches that apply end-to-end methods to directly map sensory visual inputs to actions that control the robot (e.g. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]). VSN approaches, instead of planning a geometric route over a map, learn navigation policies that enable robots to explore unknown environments while following semantically meaningful instructions, such as “Move towards a chair” or “Find a bedroom”. This particular problem is known as the Object Navigation (ObjectNav) problem [16], where an agent must navigate, in an unknown scenario, so without any map, toward an instance of a given object category using mainly visual information. All these VSN ObjectNav systems are trained and evaluated in complex simulation platforms, such as Habitat [17, 16], AI2-THOR [18] or ProcTHOR [19], where the agents learn to navigate in interactive 3D environments and complex physics-enabled scenarios. AI2-THOR and ProcTHOR enable interaction with realistic virtual 3D environments generated through computer graphics. Habitat, on the other hand, allows agents to interact with 3D models corresponding to real-world 3D scans of scenes and houses, such as those included in the HM3D datasets [20, 21].
@@
Figure 1: Traditional VSN models are trained in simulation environments where they primarily rely on RGB images. However, due to the significant domain gap between the real world and simulated images, VSN systems often fail to navigate when deployed on real robots. To address this issue, we propose the SemNav paradigm, in which VSN systems use semantic segmentation of the environment as their primary input sensor.
@@
Despite recent progress, state-of-the-art ObjectNav models often struggle with real-world generalization due to the significant domain gap between simulation-trained policies and real-world execution [22, 23, 24]. In other words, no matter how realistic the simulation environments used during learning are, the real world presents other challenges. To begin with, the images provided by the simulation platforms are either not entirely realistic or exhibit artifacts typical of 3D scanning processes of real-world scenes. The egocentric images that a robotic platform will acquire in the real world will present higher levels of blurring due to the platform’s movement, more diverse lighting situations, and will also include changes that naturally occur in dynamic environments. This results in navigation strategies that fail to generalize beyond the training domain. To address this challenge, we propose SemNav (see figure 1), a novel ObjectNav approach that uses semantic segmentation as the main visual input. Our goal in basing the visual perception of the VSN system on semantic segmentation is twofold. First, we argue that a representation based on semantic segmentation naturally mitigates the domain adaptation problem between the real world and simulated environments. The differences between the semantic segmentations of a real-world scene and those of a rendered scene in a simulation environment such as Habitat [17] are smaller than those found in RGB images from both domains. Second, by incorporating semantic priors into the visual representation of the environment, our model learns robust navigation policies that better adapt to unseen environments, hence improving the success of the navigation task. Our approach is inspired by the intuition that human navigation heavily relies on semantic cues—understanding that a “kitchen” typically contains a “sink” or a “refrigerator” enables more efficient exploration.
@@ II Related work
Robotic autonomous navigation has long been a key focus in research, with many methods developed to tackle the challenge of maneuvering robots through complex and dynamic environments. Classical navigation approaches are mainly based on Simultaneous Localization and Mapping (SLAM) [1, 2, 3, 25, 26, 27]. SLAM-based solutions primarily tackle the navigation problem by focusing on obstacle avoidance, map creation and path-planning algorithms. However, these models face challenges in generalizing to different environments [4].
To address these limitations and capitalize on recent breakthroughs in machine learning and the availability of large-scale datasets, research has increasingly shifted toward learning-based approaches for robotic navigation. Within this category, Visual Semantic Navigation (VSN) models stand out. Unlike traditional methods that depend on geometric path planning over pre-constructed maps, VSN models adopt a fundamentally different strategy by learning navigation policies that enable robots to navigate unfamiliar environments based on semantically meaningful instructions, such as “Go to a chair” (ObjectNav problem) or “Find this image” (ImageNav task). For a comprehensive overview of recent progress and challenges in VSN, interested readers may refer to [16]. Fundamentally, VSN models learn to make navigation decisions based mainly on visual information (RGB and/or depth data) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Two main VSN training approaches include both Imitation Learning (IL) and Reinforcement Learning (RL). IL-based methods learn navigation policies from previously annotated demonstrations [5, 6]. RL-based methods can be split into two categories. The first category consists of models that follow an end-to-end RL approach [28, 9, 10, 11] via interaction with the environment. Several training strategies have been proposed for these systems, including the use of auxiliary tasks [12], object-relationship graphs [13], and the fusion of visual and auditory information [29, 30]. The second category includes works that employ modular learning approaches [8, 22, 31, 23], where the navigation process is divided into independent modules addressing different aspects of navigation.
@@
As previously argued, despite significant efforts in developing VSN models, a major limitation persists: most models are trained and evaluated predominantly in simulation environments. In fact, only a few studies have attempted to evaluate VSN solutions in real-world settings [22, 23, 24]. The results reported in these studies reveal a significant domain gap problem: while VSN models achieve high success rates in simulation, their performance deteriorates drastically when deployed in real-world environments. In this work, we propose a novel VSN model, named SemNav, designed to learn navigation policies for the ObjectNav problem using a semantic segmentation of the robot’s egocentric view. The semantic segmentation output is fed into a CNN-based visual encoder, which extracts navigational cues from the segmentation information to enhance navigation in unseen environments. To the best of our knowledge, no previous VSN work has implemented this idea as we have. Some studies leverage semantic segmentation to define navigable regions and plan routes [39, 40], while others combine this information with object detection systems [41] to enable target-driven visual navigation. Our initial hypothesis, confirmed by experimental results, is that semantic segmentation serves as a valuable source of information that allows: (1) the development of more efficient VSN models and (2) the natural mitigation of domain adaptation issues.
@@ III-A SemNav dataset
The first essential requirement for training a VSN model that utilizes semantic segmentation as its visual input is the availability of a dataset containing this type of information. In this work, we have built upon the HM3D dataset [5] to construct the SemNav dataset. Specifically, in this section, we detail the process through which we enriched HM3D to enable the generation of a semantic segmentation sensor compatible with the Habitat simulator [17]. This semantic sensor allows any agent interacting with Habitat to access an egocentric view that corresponds to the semantic segmentation of the scene currently perceived by the robot.
@@
The HM3D dataset provides a collection of 3D real-world spaces, densely annotated with semantic information, as illustrated in Figure 2. This dataset comprises over 140,000 object instance annotations distributed across 216 3D environments and approximately 3,100 rooms. The semantic data is embedded in the form of texture images mapped onto the original 3D geometry of the HM3D scenes. This information is packed into binary glTF format files, one per scene. Each object instance is assigned a unique color identifier, which maps to a corresponding textual label specifying the object’s category, stored in a separate annotation file.
@@
The separation of object texture information and semantic labels into different files leads to a limitation: the built-in sensor in Habitat for the HM3D dataset does not deliver a true semantic segmentation of the scene. For instance, chairs, as illustrated in Figure 2, are not consistently assigned the same identifier or label across different scenes.
@@
Our model, SemNav, relies on semantic segmentation to navigate effectively. To meet this requirement, we have developed an enhanced semantic segmentation sensor. It builds upon the original semantic information available in the HM3D dataset. Unlike the default information, our sensor ensures that object categories are consistently assigned unique and uniform labels across all scenes. Technically, the creation of these sensors has been automated by designing a mapping process that allows translating the information associated with textures into semantic segmentation labels that are now expanded both intra- and inter-scenes. This improvement guarantees that different instances of the same object category receive identical labels within a single scene and across multiple scenes, as illustrated in Figure 2.
@@
To provide flexibility and address varying levels of semantic granularity, we have created two semantic segmentation sensors: one with 1,630 distinct object categories, referred to as SemNav 1630, and another with a more compact set of 40 categories, called SemNav 40. The SemNav 1630 sensor was generated by leveraging the fine-grained object annotations available in HM3D [5], where each manually labeled object instance was assigned to a unique category. This allowed identically annotated objects to be grouped under the same category. However, upon conducting a detailed analysis, we identified certain limitations in these annotations, including noise from overly specific object labels and occasional misclassifications. To address these challenges and improve the robustness of the semantic segmentation, we created a second semantic segmentation sensor, SemNav 40. In this version, each of the 1,630 original annotations was manually mapped to one of the 40 broader categories defined by the NYUv2 dataset [42], a widely recognized benchmark for indoor semantic segmentation. This mapping reduced the initial set of fine-grained categories to a more semantically coherent set of 40, thereby simplifying the navigation process and minimizing the impact of noisy annotations. Both SemNav 1630 and SemNav 40 were integrated into the Habitat simulator, allowing any agent to query the simulator for true semantic segmentation corresponding to its egocentric view in the 3D environment being rendered.
@@
Figure 2: Comparison between the HM3D dataset and the SemNav 1630 and SemNav 40 datasets. In HM3D, objects like chairs are inconsistently labeled with different colors across and within scenes. In contrast, SemNav 1630 assigns a uniform color to objects of the same category while reflecting finer distinctions, such as differentiating kitchen and dining tables. SemNav 40 further simplifies the labeling by merging less critical categories, grouping both types of tables under a shared label to streamline navigation tasks. Best viewed in colour.
@@ III-B SemNav model
In this work, we also introduce the SemNav model, designed to address the VSN problem known as ObjectNav [16]. The objective of an ObjectNav navigation episode is to enable an agent to navigate in a scene $S_{i}$ from a set of available scenes $\mathcal{S}=\{S_{1},\dots,S_{n}\}$, towards an object of a specific category $c_{i}$ belonging to the category set $\mathcal{C}=\{c_{1},\dots,c_{m}\}$, starting from an initial position $p_{0}$ in the navigation environment.
For our SemNav model, we define the navigation task as follows. Given a target object class $c_{i}$, e.g., chair, our goal is to navigate to an instance of this category using only the semantic segmentation of the environment. Figure 3 shows the architecture of the proposed SemNav model. We define a VSN model in which, at each position $p_{i}$, the agent has access to an environmental observation $o_{p_{i}}$, represented as the tuple:
$o_{p_{i}}=\{o^{\text{semantic}}_{i},o^{\text{compass}}_{i},o^{\text{gps}}_{i}\}\;.$
As shown in Figure 3, the SemNav model receives as inputs: the semantic segmentation of the environment ($o^{\text{semantic}}_{i}$); the relative orientation with respect to the initial viewpoint ($o^{\text{compass}}_{i}=\Delta\alpha_{i}$); and the relative displacement from the starting position ($o^{\text{gps}}_{i}=(\Delta x_{i},\Delta y_{i},\Delta z_{i})$).
Figure 3: Proposed architecture for the SemNav model.
@@
For the SemNav model, we adopt an IL strategy based on human demonstrations to learn the parameters in $\theta$. Specifically, in our experiments, we leverage a dataset of 77k human-driven navigation trajectories from HM3D [5], which are also applicable to the novel SemNav dataset. A trajectory $\tau$ of length $T$, representing a human demonstration, is defined as a sequence of observation-action tuples:$\tau=(o_{p_{0}},a_{p_{0}},o_{p_{1}},a_{p_{1}},\dots,o_{p_{T}},a_{p_{T}})$. The set of trajectories used for training is denoted as $\mathcal{T}=\big{\{}\tau_{i}\big{\}}_{i=1}^{N}$. Following an IL approach, we optimize the parameters $\theta^{*}$ of our navigation policy by solving the following optimization problem
@@
where the navigation policy is trained by minimizing the discrepancy between the probability distribution for the actions assigned by the model to each observation $o_{p_{t}}$, and the actions of the human experts.

### v2

# SemNav: Enhancing Visual Semantic Navigation in Robotics through Semantic Segmentation

@@ Abstract
Visual Semantic Navigation (VSN) is a fundamental problem in robotics, where an agent must navigate toward a target object in an unknown environment, mainly using visual information. Most state-of-the-art VSN models are trained in simulation environments, where rendered scenes of the real world are used, at best. These approaches typically rely on raw RGB data from the virtual scenes, which limits their ability to generalize to real-world environments due to domain adaptation issues. To tackle this problem, in this work, we propose SemNav, a novel approach that leverages semantic segmentation as the main visual input representation of the environment to enhance the agent’s perception and decision-making capabilities. By explicitly incorporating this type of high-level semantic information, our model learns robust navigation policies that improve generalization across unseen environments, both in simulated and real world settings. We also introduce the SemNav dataset, a newly curated dataset designed for training semantic segmentation-aware navigation models like SemNav. Our approach is evaluated extensively in both simulated environments and with real-world robotic platforms. Experimental results demonstrate that SemNav outperforms existing state-of-the-art VSN models, achieving higher success rates in the Habitat 2.0 simulation environment, using the HM3D dataset. Furthermore, our real-world experiments highlight the effectiveness of semantic segmentation in mitigating the sim-to-real gap, making our model a promising solution for practical VSN-based robotic applications. The code and datasets are accessible at https://github.com/gramuah/semnav.
@@ I Introduction
Autonomous navigation remains a fundamental challenge in robotics, particularly in unstructured and dynamic environments. Traditional approaches, such as the ones using Simultaneous Localization and Mapping (SLAM), for example, focus on a geometric reconstruction of the environment in conjunction with path planning and obstacle avoidance techniques (e.g. [12, 30]). However, these classical models typically struggle when they need to generalize across environments and domains [3].
Visual Semantic Navigation (VSN) has emerged as an alternative paradigm that leverages advances in machine learning and the availability of large-scale data. VSN models are essentially learning-based navigation approaches that apply end-to-end methods to directly map sensory visual inputs to actions that control the robot (e.g. [27, 44, 8, 7, 40, 50, 47, 46, 16]).
@@
Figure 1: Traditional VSN models are typically trained in simulated environments, relying primarily on RGB images. However, the substantial domain gap between simulated and real-world imagery often leads to poor performance when these systems are deployed on real robots. To overcome this limitation, we introduce the SemNav paradigm, where VSN systems use semantic segmentation maps of the environment as their main sensory input. This approach significantly reduces the domain gap between training simulations and real-world scenarios, enabling models to generalize more effectively and perform reliably when tested outside of simulation.
VSN approaches, instead of planning a geometric route over a map, e.g. [39, 4, 15], directly learn navigation policies that enable robots to explore unknown environments while following semantically meaningful instructions, such as “Move towards a chair” or “Find a bedroom”. This particular problem is known as the Object Navigation (ObjectNav) problem [42], where an agent must navigate, in an unknown scenario, so without any map, toward an instance of a given object category using mainly visual information. All these VSN ObjectNav systems are trained and evaluated in complex simulation platforms, such as Habitat  [36, 42], RoboTHOR [10] or ProcTHOR [11], where the agents learn to navigate in interactive 3D environments and complex physics-enabled scenarios. AI2-THOR and ProcTHOR enable interaction with realistic virtual 3D environments generated through computer graphics. Habitat, on the other hand, allows agents to interact with 3D models corresponding to real-world 3D scans of scenes and houses, such as those included in the HM3D datasets [25, 45]. VSN-based approaches have practical real-world applications, such as the development of assistive robots capable of retrieving objects and providing support to individuals in need. These models operate in a plug-and-play manner, enabling autonomous and intelligent exploration of the environment without requiring prior map construction [32]. Furthermore, VSN systems can be used for autonomous exploration of unseen or partially known environments, facilitating tasks such as search and rescue, environment monitoring, and efficient scene understanding in dynamic or unstructured spaces.
@@
Despite recent progress, state-of-the-art ObjectNav models often struggle with real-world generalization due to the significant domain gap between simulation-trained policies and real-world execution [13, 38, 16]. In other words, no matter how realistic the simulation environments used during learning are, the real world presents other challenges. To begin with, the images provided by the simulation platforms are either not entirely realistic or exhibit artifacts typical of 3D scanning processes of real-world scenes. The egocentric images that a robotic platform will acquire in the real world will present higher levels of blurring due to the platform’s movement, more diverse lighting situations, and will also include changes that naturally occur in dynamic environments. This results in navigation strategies that fail to generalize beyond the training and simulated domain. To address this challenge, we propose SemNav (see figure 1), a novel ObjectNav approach that uses semantic segmentation as the main visual input. Our goal in basing the visual perception of the VSN system on semantic segmentation is twofold. First, we argue that a representation based on semantic segmentation naturally mitigates the domain adaptation problem between the real world and simulated environments. The differences between the semantic segmentations of a real-world scene and those of a rendered scene in a simulation environment such as Habitat [36] are smaller than those found in RGB images from both domains. Second, by incorporating semantic priors into the visual representation of the environment, our model learns robust navigation policies that better adapt to unseen environments, hence improving the success of the navigation task. Our approach is inspired by the intuition that human navigation heavily relies on semantic cues—understanding that a “kitchen” typically contains a “sink” or a “refrigerator” enables more efficient exploration.
@@ II Related work
Robotic autonomous navigation has long been a key focus in research, with many methods developed to tackle the challenge of maneuvering robots through complex and dynamic environments. Classical navigation approaches are mainly based on Simultaneous Localization and Mapping (SLAM)  [12, 41, 30, 5, 31]. SLAM-based solutions primarily tackle the navigation problem by focusing on obstacle avoidance, map creation and path-planning algorithms. However, these models face challenges in generalizing to different environments [3].
To address these limitations and capitalize on recent breakthroughs in machine learning and the availability of large-scale datasets, research has increasingly shifted toward learning-based approaches for robotic navigation. Within this category, Visual Semantic Navigation (VSN) models stand out. Unlike traditional methods that depend on geometric path planning over pre-constructed maps, VSN models adopt a fundamentally different strategy by learning navigation policies that enable robots to navigate unfamiliar environments based on semantically meaningful instructions, such as “Go to a chair” (ObjectNav problem) or “Find this image” (ImageNav task). For a comprehensive overview of recent progress and challenges in VSN, interested readers may refer to [42]. Fundamentally, VSN models learn to make navigation decisions based mainly on visual information (RGB and/or depth data) [27, 44, 8, 7, 40, 50, 47, 46]. Two main VSN training approaches include both Imitation Learning (IL) and Reinforcement Learning (RL). IL-based methods learn navigation policies from previously annotated demonstrations [27, 44]. RL-based methods can be split into two categories. The first category consists of models that follow an end-to-end RL approach [52, 40, 50] via interaction with the environment. Several training strategies have been proposed for these systems, including the use of auxiliary tasks [47], object-relationship graphs [46], and the fusion of visual and auditory information [22]. The second category includes works that employ modular learning approaches [7, 13, 21, 38], where the navigation process is divided into independent modules addressing different aspects of navigation.
@@
As previously argued, despite significant efforts in developing VSN models, a major limitation persists: most models are trained and evaluated predominantly in simulation environments. In fact, only a few studies have attempted to evaluate VSN solutions in real-world settings [13, 38, 16]. The results reported in these studies reveal a significant domain gap problem: while VSN models achieve high success rates in simulation, their performance deteriorates drastically when deployed in real-world environments. In this work, we propose a novel VSN model, named SemNav, designed to learn navigation policies for the ObjectNav problem using a semantic segmentation of the robot’s egocentric view. The semantic segmentation output is fed into a Convolutional Neural Network (CNN)-based visual encoder, which extracts navigational cues from the segmentation information to enhance navigation in unseen environments. To the best of our knowledge, no previous VSN work has implemented this idea as we have. Some studies leverage semantic segmentation to define navigable regions and plan routes [2, 1], while others combine this information with object detection systems [24] to enable target-driven visual navigation. Our initial hypothesis, confirmed by experimental results, is that semantic segmentation serves as a valuable source of information that allows: (1) the development of more efficient VSN models and (2) the natural mitigation of domain adaptation issues.
@@ III-A SemNav dataset
The first essential requirement for training a VSN model that utilizes semantic segmentation as its visual input is the availability of a dataset containing this type of information. In this work, we extend the HM3D dataset [27] to create the SemNav dataset, which introduces support for a semantic segmentation sensor. Specifically, we detail the enrichment process that enables the generation of a semantic segmentation sensor compatible with the Habitat simulator [36]. This semantic sensor allows any agent interacting with Habitat to access an egocentric view that corresponds to the semantic segmentation of the scene currently perceived by the robot.
@@
The HM3D dataset provides a collection of 3D real-world spaces, densely and manually annotated with semantic information, as illustrated in Figure 2. This dataset comprises over 140,000 object instance annotations distributed across 216 3D environments and approximately 3,100 rooms. The semantic data is embedded in the form of texture images mapped onto the original 3D geometry of the HM3D scenes. This information is packed into binary glTF format files, one per scene. Each object instance is assigned a unique color identifier, which maps to a corresponding textual label specifying the object’s category, stored in a separate annotation file.
@@
The separation of object texture information and semantic labels into different files leads to a limitation: the built-in sensor in Habitat for the HM3D dataset does not deliver a true semantic segmentation of the scene. For instance, chairs, as illustrated in Figure 2, are not consistently assigned the same identifier or label across different scenes, but treated as different textures.
@@
Our model, SemNav, relies on semantic segmentation to navigate effectively. To meet this requirement, we have developed an enhanced semantic segmentation sensor. It builds upon the original semantic information available in the HM3D dataset. Unlike the default information, our sensor ensures that object categories are consistently assigned unique and uniform labels across all scenes. Technically, the creation of this sensor has been automated by designing a mapping process that allows translating the information associated with textures into semantic segmentation labels that are now expanded both intra- and inter-scenes. This improvement guarantees that different instances of the same object category receive identical labels within a single scene and across multiple scenes, as illustrated in Figure 2. Using the semantic annotation files (one per scene), it is possible to determine the semantic category associated with each object. To generate our sensor, we used these annotations to convert the packed texture information in the 3D model files into semantic segmentation data. We assigned a unique global ID and color to each annotation and replaced the original textures with their corresponding semantic segmentation colors across all scenes and rooms. After this conversion, we integrated the semantic segmentation sensor into the simulator, allowing any agent to query the simulator for the semantic segmentation of its egocentric view in the rendered 3D environment.
@@
To provide flexibility and address varying levels of semantic granularity, we have created two semantic segmentation sensors: one with 1,630 distinct object categories, referred to as SemNav 1630, and another with a more compact set of 40 categories, called SemNav 40. The SemNav 1630 sensor was generated by leveraging the fine-grained object annotations available in HM3D [27], where each manually labeled object instance was assigned to a unique category. This allowed identically annotated objects to be grouped under the same category. However, upon conducting a detailed analysis, we identified certain limitations in these annotations, including noise from overly specific object labels and occasional misclassifications. To address these challenges and improve the robustness of the semantic segmentation, we created a second semantic segmentation sensor, SemNav 40. In this version, each of the 1,630 original annotations was manually mapped to one of the 40 broader categories defined by the NYUv2 dataset [35], a widely recognized benchmark for indoor semantic segmentation. This mapping is provided as metadata within the HM3D dataset [27], which we directly used in our mapping process, for the generation of the novel semantic segmentation sensor. This reduction from fine-grained labels to 40 coherent categories simplifies navigation and minimizes the impact of noisy annotations. Both SemNav 1630 and SemNav 40 sensors were integrated into the Habitat simulator, allowing any agent to query the simulator for a true semantic segmentation corresponding to its egocentric view in the 3D environment being rendered.
@@
Figure 2: Comparison between the HM3D dataset and the SemNav 1630 and SemNav 40 novel datasets. In HM3D, objects like chairs are inconsistently labeled with different colors across and within scenes. In contrast, SemNav 1630 assigns a uniform color to objects of the same category while reflecting finer distinctions, such as differentiating kitchen and dining tables. SemNav 40 further simplifies the labeling by merging less critical categories, grouping both types of tables under a shared label to streamline navigation tasks. Figure is best viewed in colour.
@@ III-B SemNav model
In this work, we also introduce the SemNav model, designed to address the VSN problem known as ObjectNav [42]. The objective of an ObjectNav navigation episode is to enable an agent to navigate in a scene $S_{i}$ from a set of available scenes $\mathcal{S}=\{S_{1},\dots,S_{n}\}$, towards an object of a specific category $c_{i}$ belonging to the category set $\mathcal{C}=\{c_{1},\dots,c_{m}\}$, starting from an initial position $p_{0}$ in the navigation environment. The agent must reach an instance of the target category and execute a stop action before exceeding a maximum number of discrete steps, denoted as $N_{\text{max}}$. This stop action signals that the agent has recognized the successful completion of the navigation task.
For our SemNav model, we define the navigation task as follows. Given a target object class $c_{i}$, e.g., chair, our goal is to navigate to an instance of this category using only the semantic segmentation of the environment, within a maximum of 500 discrete actions $N_{\text{max}}=500$. Figure 3 shows the architecture of the proposed SemNav model, which leverages semantic segmentation information as input to guide the navigation policy. We define a VSN model in which, at each position $p_{i}$, the agent has access to an environmental observation $o_{p_{i}}$, represented as the tuple:
$o_{p_{i}}=\{o^{\text{semantic}}_{i},o^{\text{ori}}_{i},o^{\text{pos}}_{i}\}\;.$
As shown in Figure 3, the SemNav model receives as inputs: the semantic segmentation of the environment $o^{\text{semantic}}_{i}$; the relative orientation with respect to the initial viewpoint $o^{\text{ori}}_{i}=\Delta\alpha_{i}$; the relative displacement from the starting position $o^{\text{pos}}_{i}=(\Delta x_{i},\Delta y_{i},\Delta z_{i})$; the previous action taken by the agent $a_{i-1}$; and the target object class $c_{i}$.
$o_{p_{i}}=\{\,o^{\text{semantic}}_{i},\;o^{\text{ori}}_{i},\;o^{\text{pos}}_{i},\;a_{i-1},\;c_{i}\,\}.$
Figure 3: Our proposed architecture for the SemNav model, where the main visual input is a semantic segmentation.
The output of our SemNav model is a discrete navigation action that moves the agent within the environment. Formally, we define our SemNav model as a mapping between an observation $o_{p_{i}}$ and an action $a_{p_{i}}$ within the discrete set $\mathcal{A}=\{\textsc{turn\_left},\textsc{turn\_right},\textsc{move\_forward},\allowbreak\textsc{move\_backward},\textsc{stop}\}$. This mapping is learned through a navigation policy implemented by a deep learning model, as depicted in Figure 3. Let $\pi^{\textsc{SemNav}}_{\theta}(a_{p_{i}}\mid o_{p_{i}})$ denote this navigation policy, parameterized by the weights $\theta$ of the network architecture responsible for mapping observations to a probability distribution over actions in $\mathcal{A}$.
@@
where the navigation policy is trained by minimizing the discrepancy between the probability distribution for the actions assigned by the model to each observation $o_{p_{t}}$, and the actions provided by the human experts.